OpenStudy (anonymous):

simplify i^39 is the answer i...?

OpenStudy (goformit100):

yes

OpenStudy (callisto):

No... \[i^{39} = i^{36}i^{3} = (1)(i^3) = i^2i = -i\]

OpenStudy (anonymous):

I agree with @Callisto